Collingwood Caskets generally uses a before-tax MARR of 18 percent. Vincent wants to do a detailed...

  1. Collingwood Caskets generally uses a before-tax MARR of 18 percent. Vincent wants to do a detailed calculation of the cash flows associated with a new planer for the assembly line. What would be an appropriate after-tax MARR for him to use if Collingwood Caskets pays

Q3 ANSWER:

  1. 20 percent corporate taxes? (5 points):

………………………………………………………………………….

  1. 40 percent corporate taxes? (5 points)

………………………………………………………………………….
Answer #1

Before Tax MARR = After Tax MARR / (1 - Tax Rate)

a. If the Tax Rate is 20% and the before tax MARR is 18%, then

0.18 = After Tax MARR / (1 - 0.20)

or After Tax MARR = 0.80 * 0.18 = 0.144

So the After tax MARR will be 14.4% if Collingwood Caskets pays 20% corporate taxes

b.

If the Tax Rate is 40% and the before tax MARR is 18%, then

0.18 = After Tax MARR / (1 - 0.40)

or After Tax MARR = 0.60 * 0.18 = 0.108

So the After tax MARR will be 10.8% if Collingwood Caskets pays 40% corporate taxes

If you found this helpful, please rate it so that I can have higher earnings at no extra cost to you. This will motivate me to write more.